# Assortment Optimization Under History-Dependent Effects

### v2

@@
This paper examines how to plan multi-period assortments when customer utility depends on historical assortments. We formulate this problem as a nonlinear integer programming problem and propose solution methodologies for different regimes of this problem. First, we present a sequential revenue-ordered policy and show that it is optimal when historical assortments positively affect customer utility. Second, we show that the problem is NP-hard under the presence of a negative history-dependent effect (such as a satiation effect). In this case, we propose using a lifting-based framework to reformulate the problem as a mixed-integer exponential cone program that state-of-the-art solvers can solve. Additionally, we identify an optimal cyclic policy for an asymptotic regime, and we also relate its length to the customer’s memory length. Finally, we present a case study using a catering service dataset that shows that our model demonstrates good fitness and can effectively balance variety and revenue.
@@ 1 Introduction
Many companies sell products to customers who make repeat purchases or visit over multiple periods. For example, corporate cafeterias provide workday luncheons for employees, gyms offer various classes to their members every day, and in live streaming sales, streamers plan product assortments for consecutive broadcasting sessions to fans. Many studies have explored and validated that customers’ memory of historical experiences may affect their willingness to buy or retain a product (Bowden 2009, Lemon and Verhoef 2016). For example, a customer may initially feel satisfied with a product, but this satisfaction may wane over time. On the other hand, a customer may also prefer to consistently consume the same product, which leads to sustained or increased satisfaction. How to incorporate these historical effects and offer various period-by-period product assortments is therefore a critical decision for assortment planning.
Studies have shown that consideration of customers’ history-dependent effects in pricing decisions can significantly improve revenue (e.g., Wang 2016, Chen et al. 2023c). However, history-dependent assortment planning problems are underexplored due to data scarcity and the decision complexity. Many managers continue to rely only on their experience, industry knowledge, and a rough estimate of the demand in order to manually determine assortments, and ignore the impact of historical assortments on customer behavior. With increasing collection of transactional data, companies seek to leverage these data to understand and estimate customers’ preferences to build data-driven assortment planning systems. For instance, new technology companies have emerged that revolutionize traditional cafeterias by establishing membership systems, gathering and analyzing transactional data, and offering data-driven customized lunch menus on a daily basis, which leads to increased revenue and customer engagement. For further details and references in this regard, the reader is referred to Fooda’s Cafeteria Solution [^1] https://www.fooda.com/cafeteria-solutions-corporate-dining and Zerocater’s Corporate Cafeteria Solution. [^2] https://zerocater.com/corporate-cafeterias/
@@
Planning assortments using history-dependent effects is challenging because decision makers need to account for the interplay between cross-product and cross-period effects. Cross-product effects refer to the influence of other products in an assortment on purchase decisions. For instance, customer substitution behavior is commonly modelled using discrete choice models, like the multinomial logit (MNL) model (Gallego et al. 2019). Cross-period effects refer to the temporal influences on customer utility, i.e., the utility of a product in a given period depending on its offering history. Given both types of effect, the offering decisions for multiple products over multiple periods are made simultaneously. The number of candidate assortments, therefore, scales exponentially with the number of products and the planning horizon. This can be enormous for even a practically-sized problem. Moreover, modelling customer behavior under both types of effect makes the decision more complex because, for a given period, the purchase probability is a nonlinear function—specifically, a softmax function—of the history-dependent utility functions.
Figure 1: The benefits of incorporating history-dependent effects into assortment planning
@@
Despite the complexity of assortment planning with history-dependent effects, the benefit of considering such effects is significant. Using the data from a corporate cafeteria with which we cooperated, we evaluate assortments derived from the MNL model, from a history-dependent model (our model [^3] We estimate the history-dependent effect and memory length (i.e., two periods) using transactional data of a corporate cafeteria. Please refer to Section 5 for details.), and the current practice observed in the data (as a benchmark to normalize the revenue). The results shown in Figure 1 indicate that our model achieves the highest average daily revenue in out-of-sample test, approximately 19% more than the revenue generated by the MNL model. Our model can also produce assortments that have greater variety than those produced by the MNL model. The current practice decided by managers resulted in assortments with a greater variety but also the lowest revenue. See Section 5 for detailed discussion on the case study.
@@ 1.1 Description of Results
We study a multi-period assortment planning problem in which a firm facing both cross-product and cross-period effects needs to decide what to offer in each period. We model customer choice in a given period using a variant of the MNL model in which the utility of each product linearly depends on its past offering decisions. In our model, an assortment decision for a given period affects both current and future purchase behavior in a nonlinear fashion. In other words, our assortment planning problem is a nonlinear integer programming problem.
We propose solution methodologies for different regimes of this problem. First, we present a sequential revenue-ordered (RO) policy and show that it is optimal if historical assortments positively affect customer utility, i.e., there is an addiction effect. This characterization indicates that our problem with the addiction effect is solvable in polynomial time. Second, we prove that when historical assortments negatively affect the current utility, i.e., there is a satiation effect, the problem becomes NP-hard even when there are only two planning horizons. In this case, we propose a mixed-integer exponential cone program (MIECP) formulation to find a global optimal solution. The key idea behind our formulation is to lift the original nonconvex constraints into a higher-dimensional space where the lifted constraints are simpler to model. This simpler reformulation allows us to utilize prevalent techniques in the optimization literature, such as the McCormick envelope, the perspective formulation, and the Lovász extension, to model the lifted structures. Numerical studies indicate that our formulation efficiently achieves optimality and yields higher revenue than heuristic policies, especially under strong satiation effects. For problems with a long-term planning horizon, theoretically, we show that a cyclic policy is optimal, and find that, under a certain non-overlapping condition, the optimal cycle length equals to the customer memory length plus one. Computationally, finding such cyclic policies is NP-hard but also admits an MIECP formulation. Moreover, for the case with a memory length of one, we derive a mixed-integer linear programming (MILP) formulation that is stronger than the conic formulation.
We conduct a case study using transactional data from a corporation cafeteria operated by a high-tech catering service company in China. The cafeteria primarily serves workday lunches for employees who work nearby. We estimate customers’ utility functions and then design weekly day-by-day lunch menus. The estimation shows that a satiation effect does exist, but it only lasts for a short period (i.e., two or three days). We demonstrate that our history-dependent choice model can provide menus that balance variety and revenue, while the menus used in practice sacrifice revenue and the menus proposed by the classic MNL model have low variety and low revenue.
@@
    Our model, which we refer to as HAP, simultaneously incorporates the cross-period and cross-product effects into the multi-period assortment planning process. We demonstrate via a case study of a corporate cafeteria that the operations in our model have several benefits. First, our result suggests that the service provider should provide a greater variety of products if satiation effects exist. When satiation effects last longer, it is beneficial to offer more types of products across periods. Furthermore, it is recommended that a product be offered only once within the memory length in order to prevent a utility decrease due to satiation effects, particularly in the case of products with high prices. This observation naturally leads to a “non-overlapping” cyclic assortment planning policy, which we explore in Section 4, where we establish its optimality. Most importantly, our history-dependent policy effectively balances variety and the revenue objective, resulting in assortments that achieve the highest revenue in out-of-sample tests while maintaining the same variety as found in practical scenarios.
@@
    We characterize, in terms of history-dependent effects, the conditions under which a sequential revenue-ordered policy globally solves our HAP model. In general, it can be shown that finding a global optimal solution is NP-hard. To tackle this difficulty, we develop an MIECP formulation of our HAP model by using lifting and coupling ideas from the optimization literature, including perspective formulation, the McCormick envelope, and the Lovász extension. Computational studies show that our formulation and its variant outperform a MILP formulation of (HAP) obtained using a prevalent strategy in the mixed-integer nonlinear programming (MINLP) literature and the original nonlinear integer formulation (HAP) solved with the state-of-the-art MINLP solver SCIP.
@@
    We establish the asymptotical optimality of cyclic policies, and finding such policies allows an MIECP formulation. Our study reveals a connection between the length of an optimal cyclic policy and the length of customer memory. In particular, we characterize the condition under which the optimal cyclic length is the customer memory length plus one. This condition, in addition, leads to an MILP formulation for finding an optimal two cyclic policy, which is, theoretically, tighter than the previous conic formulation.
@@ 1.2 Literature Review
Extensive research has incorporated history-dependent effects into operations problems, including pricing and promotion (Wang 2016, Bi et al. 2020, Chen et al. 2023c), service design (Bernstein et al. 2022, Lei et al. 2023, Fan et al. 2024), resource allocation (Adelman and Mersereau 2013), retention management (Bastani et al. 2022, Kanoria et al. 2024), etc. The research shows that ignoring history-dependent effects in customer behavior may lead to substantial losses. For instance, considering a reference price in choice estimation can reduce a $5.82\%$ prediction error (Wang 2018), and the optimal bundle accounting for customer satiation obtains a revenue lift greater than $4.5\%$ (Gürlek et al. 2023).
History-dependent effects are receiving increasing attention in assortment planning. For instance, Chen et al. (2023b) study personalized assortment planning problems when customers’ choice probabilities depend on historical purchases. Several recent studies focus on designing online assortment algorithms to learn customers’ satiation levels and maintain high engagement (e.g., Cao et al. 2019, Bastani et al. 2022, Sumida and Zhou 2023). Our study investigates the choice-based offline assortment planning problem, such as the daily menu design for corporate cafeterias. Unlike online recommendation settings, the offline assortment planning problem determines multi-period assortments at the beginning of a planning horizon.
@@
A prevalent strategy for solving assortment optimization is to construct revenue-ordered assortments. Talluri and Van Ryzin (2004) show that the optimal solution of the assortment optimization problem under the MNL model is revenue-ordered. Wang and Wang (2017) propose a quasi-revenue-ordered policy to solve assortment optimization with endogenous network effects. Gao et al. (2021) prove that the revenue-ordered structure is optimal for the assortment optimization problem under the MNL model when the offered assortment is gradually revealed. Xu and Wang (2023) exploit the revenue-ordered property to solve a multistage assortment optimization problem. We contribute to this literature by showing that a sequential revenue-ordered policy is optimal for our problem under the addiction effect.
@@
Last, we comment that our paper contributes to a growing set of applications of conic programming/discrete optimization in designing operations systems, including joint inventory-location problems (Atamtürk et al. 2012), designing battery swap networks for electric vehicles (Mak et al. 2013), scheduling (Kong et al. 2013), process flexibility design (Yan et al. 2018, Simchi-Levi and Wei 2012), electric vehicle charging (Chen et al. 2023a), proactive policing (He et al. 2023a), network flow (Simchi-Levi et al. 2019).
@@ 1.3 Structure
The rest of the paper is organized as follows. In Section 2, we formally define the assortment planning problem, show its NP-hardness, and identify a case that is solvable in polynomial time. In Section 3, we propose an equivalent MIECP formulation. In Section 4, we present our results on cyclic policies. We present a case study using real data in Section 5 and report numerical results on synthetic data in Section 6. We conclude the paper in Section 7. All proofs are given in the Appendix.
@@ 2 Model
We define our multi-period assortment planning problem with history-dependent effects in Section 2.1. In Section 2.2, we then show the NP-hardness of our HAP model when history-dependent effects are negative. In Section 2.3, we identify a case that is solvable in polynomial time. Specifically, we show in Theorem 2.2 that a sequential revenue-ordered policy is optimal if the history-dependent effects are positive.
2.1 Assortment Planning under History-Dependent Effects
@@
Consider a firm that sells $N$ products to customers. Let $[N]:=\{1,\ldots,N\}$ [^4] In this paper, $[k]:=\{1,2,\dots,k\}$ for integer $k\in\operatorname{\mathbb{Z}}_{+}$. denote the set of all products, and let $[N]^{+}:=[N]\cup\{0\}$, where we use the index 00 to denote the outside option. For each product $i\in[N]$, let $r_{i}$ denote the revenue obtained from selling an instance of product $i$. We assume that products are indexed such that $r_{1}\geq r_{2}\geq\dots\geq r_{N}>0$. There is a total of $T$ planning horizons, and the firm needs to decide what to offer in each period. More specifically, for each period $t\in[T]$, let $\boldsymbol{x}^{t}=(x^{t}_{1},\ldots,x^{t}_{N})\in\{0,1\}^{N}$ be a binary decision variable modelling the subset of products offered in period $t$, i.e., $x^{t}_{i}=1$ if and only if product $i$ is offered in period $t$. We assume that in each period, one unit mass of customers arrives, and the customers select an item from the offered products or leave without a purchase. For example, in a setting such as corporate cafeterias, the firm serves relatively stable customers (e.g., most customers are employees who work nearby), and customers make purchase decisions on a daily basis.
To model the customer choice behavior in each period, we modify the MNL model to encapsulate the impact of the historical assortments on the current choice behavior. Under the random utility model (Gallego et al. 2019, Section 4), a customer assigns a random utility to each product and selects the product that yields the highest utility. When modeling the utility of each product, we take into account information about the history of assortments. More specifically, the utility of product $i$ in period $t$ affinely depends on its historical offering records within a memory length $M$:
$\beta_{i}^{0}+\sum_{m\in[M]}x^{t-m}_{i}\beta_{i}^{m}+\xi_{i}^{t},$ (Utility)
@@
where $M\in\operatorname{\mathbb{Z}}_{+}$ denotes the customer memory length and $\xi_{i}^{t}$ is a random variable. For simplicity in the presentation, we assume throughout this paper that $x_{i}^{k}:=0$ for $k\leq 0$. The parameter $\beta^{0}_{i}$ is the base utility of the product $i$ that can be determined by various factors such as brand, flavor, price, and so on. The parameter $\beta_{i}^{m}$ is the incremental utility if product $i$ is offered in the $m^{\text{th}}$ period before the current time. [^5] In our planning models, the assortments are fixed for each period. Therefore, we do not consider personalized assortments or model individual-level utility functions. Let $\boldsymbol{\beta}:=(\beta_{i}^{m})_{i\in[N],m\in[M]}$ denote the history-dependent effect matrix. Note that the value of $\beta_{i}^{m}$ can be either positive or negative. If $\beta_{i}^{m}$ is positive, exposure to product $i$ increases a customer’s preference for it and thus functions as a stimulus (Fox et al. 1997). Conversely, if $\beta_{i}^{m}$ is negative, product $i$ has a satiation effect, which leads to customer fatigue from repeated offerings.
@@
As in the MNL model, we assume that random variables $\xi_{i}^{t}$ are i.i.d. as the standard Gumbel distribution. For a given assortment planning $\boldsymbol{x}:=(\boldsymbol{x}^{1},\cdots,\boldsymbol{x}^{T})$, we use $\pi^{t}_{i}(\boldsymbol{x}^{t},\ldots,\boldsymbol{x}^{t-M})$ to denote the probability of purchasing product $i$ in period $t$. Under the MNL model, the probability of purchasing product $i$ in period $t$ is
@@
$\pi^{t}_{0}(\boldsymbol{x}^{t},\ldots,\boldsymbol{x}^{t-M})$ $=\dfrac{1}{1+\sum_{j\in[N]}x^{t}_{j}\exp\bigl{(}\beta_{j}^{0}+\sum_{m\in[M]}x^{t-m}_{j}\beta^{m}_{j}\bigr{)}}.$
The purchase probabilities depend on two components: the current assortment $\boldsymbol{x}^{t}$ and the past assortments $\boldsymbol{x}^{t-1},\dots,\boldsymbol{x}^{t-M}$. The former dependence captures cross-product effects such as the substitution effect among products, while the latter dependence models cross-period effects.
With these definitions, assortment planning under history-dependent effects can be expressed as the following fractional binary program, called the history-dependent assortment planning problem (HAP):
@@
where $\mathcal{X}$ (resp. $\mathcal{P}$) is a system of linear inequalities that models cross-product constraints (resp. cross-period constraints). The cross-product constraints often include resource restrictions within a single period, such as cardinality and capacity constraints, and address various operational considerations in the assortment planning literature (Sumida et al. 2021, Chen et al. 2024). Cross-period constraints describe the intertemporal relationships of assortments across different periods, such as non-overlapping assortments over consecutive periods, which naturally occur in multi-period assortment planning problems (Liu et al. 2020, Chen et al. 2023b).
2.2 Hardness Result
This subsection shows that our model (HAP) is NP-hard to solve when history-dependent effects are negative. In particular, when products have identical revenue and the planning horizon is two, we can reduce a NP-hard problem, the 3/4-Partition problem, to our problem (HAP).

### v3

@@
This paper examines how to plan multi-period assortments when customer utility depends on historical assortments. We formulate this problem as a nonlinear integer programming model and show it is NP-hard in the presence of a negative history-dependent effect (such as a satiation effect). We build solution methodologies for obtaining global optimal solutions under a general setting that the history-dependent effects could be a mixture of positive and negative. We propose using a lifting-based framework to reformulate the problem as a mixed-integer exponential cone program that state-of-the-art solvers can solve. We also design a sequential revenue-ordered policy and show that it solves our problem to optimality in polynomial time when historical assortments positively affect customer utility (such as an addiction effect). Additionally, we identify an optimal cyclic policy for an asymptotic regime, and we also relate its length to the customer’s memory length. Finally, we present a case study using a catering service dataset, showing that our model demonstrates good fitness and can effectively balance variety and revenue.
@@ 1 Introduction
Many companies sell products to customers who make repeat purchases or visit over multiple periods. For example, corporate cafeterias provide workday luncheons for employees; online grocery flash sell platform plans assortments for customers who repeatedly visit the online store and seek deep discounts. Customers’ utility may vary over time and may divert to products offered by competitors if they feel “bored”. Therefore, these companies must carefully plan their daily assortments over time to keep their customers satisfied.
A typical example is the corporate dining industry. Corporate dining is an essential but often overlooked industry. It was reported that a working person will spend an average of \$2500 per year on working-day lunches (ezCater 2024). Though corporate cafeterias are convenient and usually offer subsidized prices, they often face complaints of boredom due to the need for more dish variety. To address this, cafeteria managers manually diversify their menus based on their experience and industry knowledge. However, without a careful understanding of how previous menus impact customer demand, manual changes may be inefficient.
Many studies have explored and validated that customers’ historical experiences may affect their willingness to buy or retain a product (Bowden 2009, Lemon and Verhoef 2016). For example, a customer may initially feel satisfied with a product, but this satisfaction may wane over time. This utility decay phenomenon is also observed in fashion products (Caro et al. 2014). On the other hand, a customer may also prefer to consistently consume “traditional and familiar” food that could maintain or increase her utility (Edenred 2024). Today, with the increasing collection of transactional data, companies seek to leverage these data to estimate customers’ utility changes and build data-driven assortment planning systems. For instance, new technology companies have emerged that revolutionize traditional catering services by analyzing transactional data to estimate customers’ preferences and offering optimized daily cafeteria menus (Fooda 2024).
Besides corporate cafeterias, analogous situations arise in various other sectors. Flash sales platforms must organize their promotional assortments ahead of time, based on the products suppliers provide (de Albéniz et al. 2020). For instance, Meituan, a leading Chinese online shopping platform, manages a grocery flash sale channel. Due to the lead time of contract negotiations with suppliers, Meituan needs to decide, in advance, assortments for daily deals over a fixed planning horizon. Given that customers frequently visit the channel for discounted groceries, planning must consider historical offerings. More details about Meituan’s flash sale campaign are in the ecompanion EC.1. Managers of cafeterias and flash sales encounter a common multi-period planning problem that decides how to offer identical assortments to all customers over a fixed planning horizon. In practice, these assortments are usually predetermined before the planning horizon, possibly due to limits (such as lead time) from the supplier side.
@@
In this paper, we focus on the above applications and investigate the corresponding multi-period assortment planning problem considering history-dependent utilities to maximize the long-run average revenue. The history-dependent effects bring new challenges for assortment planning because of the interplay between cross-product and cross-period effects. Cross-product effects refer to the influence of other products in an assortment on purchase decisions. For instance, customer substitution behavior is commonly modeled using discrete choice models, like the multinomial logit (MNL) model (McFadden 1974). Cross-period effects refer to the temporal influences on customer utility, i.e., the utility of a product in a given period depending on its offering history. Given both types of effects, the offering decisions for multiple products over multiple periods are made simultaneously. The number of candidate assortments, therefore, scales exponentially with the number of products and the planning horizon. This can be enormous for even a practically-sized problem. Moreover, customer behavior under both types of effect makes the decision more complex because, for a given period, the purchase probability is a nonlinear function—specifically, a softmax function—of the history-dependent utility functions.
[…]
The average out-of-sample results of three assortment plans Our model is the history-dependent model. We normalize the revenue generated by the current practice to one. The assortment size is the number of product types offered in the planning horizon.
@@
Despite the complexity of assortment planning with history-dependent effects, the benefit of considering such effects is significant. Using the data from a corporate cafeteria with which we cooperated, we evaluate assortments derived from a history-dependent model (our model, estimated in Section 5), from the MNL model, and the current practice observed in the data (as unit one for normalization). The results shown in Figure 1 indicate that our model achieves the highest revenue in the out-of-sample test, about $10.4\%$ more than the revenue generated by the MNL model. Our model also produces assortments that have greater variety than those obtained by the MNL model. The current practice decided by managers resulted in assortments with a greater variety but the lowest revenue. See Section 5 for details of the case study.
@@ 1.1 Description of Results
We study a multi-period assortment planning problem in which a firm facing both cross-product and cross-period effects needs to decide what to offer in each period to maximize the average expected revenue. We model customer choice in a given period using a variant of the MNL model in which the utility of each product linearly depends on its past offering decisions. In our model, an assortment decision for a given period affects both current and future purchase behavior in a nonlinear fashion. In other words, our assortment planning problem is a nonlinear integer programming problem.
We show that this problem is NP-hard under the presence of negative history-dependent effects and propose exact solution methodologies for a general setting that the history-dependent effects could be a mixture of addiction and satiation. We propose a lifting-based reformulation framework to obtain a mixed-integer exponential cone program (MIECP) formulation. The key idea behind our formulation is to lift the original nonconvex constraints into a higher-dimensional space where the lifted constraints are simpler to model. This reformulation allows us to utilize prevalent techniques in the optimization literature, such as the McCormick envelope, the perspective formulation, and the Lovász extension, to model the lifted structures. Numerical studies indicate that our formulation efficiently achieves optimality and yields higher revenue than heuristic policies. We also identify an optimal sequential revenue-ordered (RO) policy if history-effects are all positive, i.e., addiction effects. This characterization indicates that our problem with the addiction effect is solvable in polynomial time. For problems with a long-term planning horizon, theoretically, we show that a cyclic policy is optimal and find that, under a certain non-overlapping condition, the optimal cycle length equals the customer memory length plus one. Computationally, finding such policies is NP-hard but also admits an MIECP formulation. Moreover, the non-overlapping condition leads to a mixed-integer linear programming (MILP) formulation that is stronger than the conic one.
We conduct a case study using transactional data from a corporation cafeteria operated by a high-tech catering service company in China. The cafeteria primarily serves workday lunches for employees who work nearby. We estimate customers’ utility function and then design weekly day-by-day lunch menus. The estimation shows that a satiation effect does exist, but it only lasts for a short period (i.e., two or three days). We demonstrate that our model can provide menus that balance variety and revenue, while current industry practices sacrifice revenue, and the classic MNL model yields menus with limited variety and low revenue.
@@
    Our model, which we refer to as HAP (i.e., history-dependent assortment problem), simultaneously incorporates the cross-product and cross-period effects into the multi-period assortment planning process. The cross-period history-dependent effects could be all positive, all negative, and mixed across products and periods. We demonstrate, via a case study of a corporate cafeteria, that our model has several benefits. First, our result suggests that the service provider should provide a greater variety of products if satiation effects exist. When satiation effects last longer, it is beneficial to offer more types of products across periods. Furthermore, it is recommended that a product be offered only once within the memory length to prevent a utility decrease due to satiation effects, particularly in the case of products with high prices. This observation naturally leads to a “non-overlapping” cyclic assortment planning policy, which we explore in Section 4, where we establish its optimality. Most importantly, our history-dependent policy effectively balances variety and the revenue objective, resulting in assortments that achieve the highest revenue in out-of-sample tests while maintaining the same variety as found in practical scenarios.
@@
    We characterize, in terms of history-dependent effects, the conditions under which a sequential revenue-ordered policy globally solves our HAP model. In general, it can be shown that finding a global optimal solution is NP-hard. To tackle this difficulty, we develop an MIECP formulation of our HAP model by using lifting and coupling ideas from the optimization literature, including perspective formulation, the McCormick envelope, and the Lovász extension. In particular, we modify the Lovász extension to model attraction values under general history-dependent effects. Our formulations, together with cutting-plane algorithms, outperform an MILP formulation of HAP obtained using a prevalent strategy in the mixed-integer nonlinear programming (MINLP) literature and the original nonlinear integer formulation HAP solved by the state-of-the-art MINLP solver SCIP.
@@
    We establish the asymptotic optimality of cyclic policies, and finding such policies allows an MIECP formulation. Our study reveals a connection between the length of an optimal cyclic policy and the length of customer memory. In particular, we characterize the condition under which the optimal cyclic length is the customer memory length plus one. This condition, in addition, leads to an MILP formulation for finding optimal cyclic policies, which is, theoretically, tighter than the previous conic formulation.
@@ 1.2 Literature Review
Extensive research has incorporated history-dependent effects into operations problems, including pricing and promotion (Bi et al. 2025, Chen et al. 2023c), service design (Bernstein et al. 2022, Lei et al. 2023), resource allocation (Adelman and Mersereau 2013), retention management (Kanoria et al. 2024), etc. Studies show that ignoring history-dependent effects may lead to substantial losses. For instance, considering the reference price can reduce $5.82\%$ prediction error of sales (Wang 2018), and the optimal bundle accounting for customer satiation improves the revenue by more than $4.5\%$ (Gürlek et al. 2023).
History-dependent effects are receiving increasing attention in assortment planning. One stream of work is to design personalized assortment planning for online platforms. The history-dependent effect derives from repeated interactions between the firm and customers, which breaks the traditional single-interaction assumption in personalized assortment optimization literature (El Housni and Topaloglu 2023). Based on the easily accessible purchase history of a customer, these studies consider two settings: (i) customers have stable repeated interactions with the firm, e.g., Chen et al. (2023b) optimize adaptive assortment decisions based on information inferred from a customer’s historical purchases; (ii) a customer can leave or churn if he or she is not satisfied. Hence, many studies focus on designing delicate exploration-exploitation recommendation algorithms that may limit exploration budgets (Bastani et al. 2022, Sumida and Zhou 2023).
For non-personalized assortment planning, the history-dependent effect is captured by a dynamic choice/demand function of the entire market. The effect may derive from attributes of a product itself, e.g., declining utility. Such intertemporal connections lead to a nonlinear and high-dimensional problem. Two settings are considered: (i) an adaptive assortment policy, e.g., Caro and Martínez-de Albéniz (2020) consider a content release problem where the attraction of content decays and the number of followers stochastically evolves; (ii) a non-adaptive assortment plan for the entire horizon made before the first period, e.g., Caro et al. (2014) optimize release schedule of fast fashion products with exponentially declining attraction before period one due to a long lead time. Our work contributes to the second type of literature. Our model can represent the model in Caro et al. (2014) and adds more flexibility (see details in Section 2.2), which brings new challenges and needs additional methods.
@@
Our paper contributes to the growing literature on developing new solution methods for assortment planning. The three primary approaches include the revenue-ordered policy, approximation algorithms, and integer programming for obtaining exact solutions. Since Talluri and Van Ryzin (2004) showed that the optimal assortment under the MNL model is revenue-ordered, this policy has been extensively studied in various contexts (Wang and Wang 2017, Sumida and Zhou 2023). For example, Gao et al. (2021) prove that the revenue-ordered assortment policy is optimal under the MNL model when the offered assortment is gradually revealed. Xu and Wang (2023) exploit the revenue-ordered property to solve a multistage assortment optimization problem. We contribute to this literature by showing that a sequential revenue-ordered policy is optimal for our problem under the addiction effect.
A stream of studies focuses on designing approximation algorithms for assortment plannings. For instance, Feldman and Topaloglu (2015) and Désir et al. (2020) design constant factor approximation algorithms for capacitated assortment optimization under the nested logit model and the Markov chain choice model, respectively. Later, Désir et al. (2022) propose a fully polynomial time approximation scheme (FPTAS) for capacitated assortment problems under the mixed MNL, Markov chain, and nested logit choice models. FPTAS are also proposed to solve assortment optimization under the exponential choice model (Aouad et al. 2023), with online dynamic behavior (Liu et al. 2020, Feldman and Jiang 2023, Aouad et al. 2024), and with multi-purchase behavior (Jasin et al. 2024, Luan et al. 2025).
@@
Last, we comment that our paper contributes to a growing set of applications of conic programming/discrete optimization in designing operations systems, including joint inventory-location problems (Atamtürk et al. 2012), battery swap networks design (Mak et al. 2013), service delivery scheduling (Kong et al. 2013), process flexibility design (Simchi-Levi and Wei 2012, Yan et al. 2018), electric vehicle charging (Chen et al. 2023a), proactive policing (He et al. 2023), network flow (Simchi-Levi et al. 2019), and nonlinear resource allocation (He and Tawarmalani 2024).
@@ 1.3 Structure
The rest of the paper is organized as follows. In Section 2, we formally define the assortment planning problem and establish its computational complexity. In Section 3, we propose a mixed-integer exponential cone formulation that can handle any mixed addiction-satiation history-dependent effects. We prove that, under positive effects, a revenue-ordered policy is optimal. In Section 4, we present our results on cyclic policies. We present a case study using real data in Section 5 and report numerical results on synthetic data in Section 6. We conclude the paper in Section 7. All proofs are given in the ecompanion.
@@ 2 Model
In Section 2.1, we define our multi-period assortment planning problem with history-dependent effects and establish its computational complexity. In Section 2.2, we discuss the model features in detail.
2.1 Incoporating History-Dependent Effects into Choice Models
@@
Consider a firm that sells $N$ products to customers. Let $[N]:=\{1,\ldots,N\}$ denote the set of all products, and let $[N]^{+}:=[N]\cup\{0\}$, where the index 00 denotes the outside option. For each product $i\in[N]$, let $r_{i}$ denote the revenue obtained from selling an instance of product $i$. There is a total of $T$ planning horizons, and the firm needs to decide what to offer in each period. More specifically, for each period $t\in[T]$, let $\boldsymbol{x}^{t}=(x^{t}_{1},\ldots,x^{t}_{N})\in\{0,1\}^{N}$ be a binary decision variable modeling the subset of products offered in period $t$, i.e., $x^{t}_{i}=1$ if and only if product $i$ is offered in period $t$. We assume that in each period, one unit mass of customers arrives, and the customers select an item from the offered products or leave without a purchase. For example, in a setting such as corporate cafeterias, the firm serves relatively stable customers (e.g., most customers are employees who work nearby), and customers make purchase decisions on a daily basis.
We modify the MNL model to encapsulate the impact of the historical assortments on customer choices. Each product has a utility $U_{i}$, which depends on attributes of the product and environment. Following Guadagni and Little (1983), Wang (2018) and Xie and Wang (2024), we assume that the utility of one product is a linear function of its attributes and historical offerings. Specifically, the utility of product $i$ in period $t$ affinely depends on its historical offering records within a memory length $M$:
$U_{i}^{t}=\beta_{i}^{0}+\sum_{m\in[M]}x^{t-m}_{i}\beta_{i}^{m},$ (Utility)
@@
where $M\in\operatorname{\mathbb{Z}}_{+}$ denotes the customer memory length. For simplicity in the presentation, we assume throughout this paper that $x_{i}^{k}:=0$ for $k\leq 0$. The parameter $\beta^{0}_{i}$ is the base utility of the product $i$ determined by its attributes. The parameter $\beta_{i}^{m}$ is the incremental utility if product $i$ is offered in the $m^{\text{th}}$ period before the current time. Let $\boldsymbol{\beta}:=(\beta_{i}^{m})_{i\in[N],m\in[M]}$ denote the history-dependent effect matrix. The value of $\beta_{i}^{m}$ can be positive or negative. A positive $\beta_{i}^{m}$ indicates that exposure to product $i$ increases its utility, acting as a stimulus (Fox et al. 1997). While a negative $\beta_{i}^{m}$ suggests a satiation effect, leading to customer fatigue from repeated offerings. The history-dependent effects can vary across products and periods. In our planning model, the assortments are fixed for each period. Therefore, we do not consider personalized assortments or model individual-level utility functions.
@@
For a given assortment plan $\boldsymbol{x}:=(\boldsymbol{x}^{1},\cdots,\boldsymbol{x}^{T})$, we use $\pi^{t}_{i}(\boldsymbol{x}^{t},\ldots,\boldsymbol{x}^{t-M})$ to denote the purchase probability of product $i$ in period $t$. Under the MNL model, the purchase probability of product $i$ in period $t$ is
@@
$\pi^{t}_{0}(\boldsymbol{x}^{t},\ldots,\boldsymbol{x}^{t-M})$ $=\dfrac{1}{1+\sum_{j\in[N]}x^{t}_{j}\exp\bigl{(}\beta_{j}^{0}+\sum_{m\in[M]}x^{t-m}_{j}\beta^{m}_{j}\bigr{)}},$
where we normalize the utility of the outside option to 0. The purchase probabilities depend on two components: the current assortment $\boldsymbol{x}^{t}$ and the past assortments $\boldsymbol{x}^{t-1},\dots,\boldsymbol{x}^{t-M}$. The former dependence captures cross-product effects such as the substitution effect, while the latter dependence models cross-period effects. The purchase probability can also be viewed as the market share of product $i$ in period $t$ and presents the effect of historical assortments on sales of the entire market (Batsell and Polking 1985).
With these definitions, assortment planning under history-dependent effects can be expressed as the following fractional binary program, called the history-dependent assortment problem:
@@
where $\mathcal{X}$ (resp. $\mathcal{P}$) is a system of linear inequalities that models cross-product constraints (resp. cross-period constraints). The cross-product constraints often include resource restrictions within a single period, such as cardinality and capacity constraints, and address various operational considerations in the assortment planning literature (Sen et al. 2018, Chen et al. 2024). Cross-period constraints describe the intertemporal relationships of assortments across different periods, such as non-overlapping assortments over consecutive periods, which occur in multi-period assortment planning problems (Liu et al. 2020, Chen et al. 2023b).
This model is generally difficult to solve. We show in the following proposition that  (HAP) is NP-hard under very simple settings if history-dependent effects are negative.